1. Build up the operation tables for group G with orders 1, 2, 3 and 4 using the elements a, b, c, and e as the identity element in an appropriate way.
2. i. State the Lagrange’s theorem of group theory.
ii. For a subgroup H of a group G, prove the Lagrange’s theorem.
iii. Discuss whether a group H with order 6 can be a subgroup of a group with order 13 or not. Clearly state the reasons.
1
Expert's answer
2020-06-11T20:48:31-0400
  1. We denote by "*" an operation of the group "G" . We point out that the group can not have elements of order 3. I.e., if "b^3=e" then the operation is wrongly defined. We point out that the general number of possible definitions of operation "*" on the group "G" is less than 9. As an example, "*" can be defined in the following way:


"\begin{matrix}
 * && e & a& b& c\\
\\
 e &&e & a& b& c\\
 a& & a & e&c& b\\
b && b & c& a& e\\
c && c& b& e& a\\
\end{matrix}"

The latter table is for group of order 4. We point out that if we consider only first row and first column, we obtain a table for the group of order 1 and if we remain first two rows and columns we get a table of order 2. It is not possible to obtain a table for a group of order 3 from the table for the group of order 4 (see Lagrange’s theorem). Therefore we present a table for a group of third order separately:

"\begin{matrix}
 * && e & a& b\\
\\
 e &&e & a& b\\
 a& & a & b&e\\
b && b & e& a\\
\end{matrix}"


2.(i) Lagrange’s theorem.

For every finite group "G" the order of every subgroup "H" divides the order of "G" .


(ii) It is enough to prove that groups of order 3 do not exist. Suppose "H" is a subgroup of order 3 with elements "e,a,b" . Let us try to extend the group operation on element "c" . Suppose that "c*a=x" , where x is from the set of elements "e, a, b, c." We multiply by "x^{-1}" and receive that "c=x*a^{-1}" . Thus, "c" has to belong to "H" . It shows that it is not possible to extend the group property from 3 to 4 elements. It completes the proof.


(iii) Due to the Lagrange's theorem, a group of order 6 can not be a subgroup, since 6 does not divide 13.